OpenStudy (anonymous):

How do you find the exact value for csc (7pi/6) and cot (-pi/3)

zepdrix (zepdrix):

\(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf \cot(-\pi/3)\quad=\quad \frac{1}{\tan(-\pi/3)}\]Do you know how to find tan(-pi/3)?

OpenStudy (ksaimouli):

Csc= 1/sin

OpenStudy (jdoe0001):

use the Unit Circle :)

zepdrix (zepdrix):

Tangent is an `odd` function, So this holds true,\[\Large\bf\sf \tan(-\theta)\quad=\quad -\tan(\theta)\] So we can further simplify it to,\[\Large\bf\sf \cot(-\pi/3)\quad=\quad \frac{1}{\tan(-\pi/3)}\quad=\quad \frac{1}{-\tan(\pi/3)}\]And then yah, use your unit circle from there :o

OpenStudy (anonymous):

would csc's final answer be -2?

OpenStudy (ksaimouli):

Yup

OpenStudy (anonymous):

and would cot's answer be \[\sqrt{3}\]?

OpenStudy (jdoe0001):

\(\bf tan(\theta)=\cfrac{{\color{red}{ \square }}}{{\color{blue}{ \square }}}\qquad cot(\theta)=\cfrac{{\color{blue}{ \square }}}{{\color{red}{ \square }}} \\ \quad \\ \quad \\ tan\left(\cfrac{-\pi}{3}\right)=\cfrac{{\color{red}{ \square }}}{{\color{blue}{ \square }}}\qquad cot\left(\cfrac{-\pi}{3}\right)=?\)